Steve is half as old as Stacey. The sum of their ages five years ago was 8. How old are Steve and Stacey now?
I got the answer...
Steve is 7 and Stacey is 10

Answer by alicealc(293)

Steve's age = V
Stacy's age = C

V = 1/2 * C
2V = C

five years ago:
V - 5 + C - 5 = 8
V - 5 + 2V - 5 = 8
3V - 10 = 8
3V = 8 + 10
3V = 18
V = 18/3 = 6

So, Steve is 6 years old and Stacy is 2*6 = 12 years old